OpenStudy (anonymous):

HELP PLEASE! Find Y if the distance between a and b is 25and b is in the 1st quadrant A=(3,-18) B=(10,y)

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

use distance formula as written below.. \[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\] here d=25, x1=3 x2=10 y1=-18 y2=y put values and solve for y..

OpenStudy (anonymous):

What is confusing me is (y-18)^2 would it be y^2+324+49?

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

(y-18)^2=y2-36y+324

OpenStudy (anonymous):

actually its (y+18) ^2 = y^2+36y+324

OpenStudy (anonymous):

tell us what Y you got when you finished solving the problem.

OpenStudy (anonymous):

I got -6 for Y @galacticwavesXX @zaibali.qasmi

OpenStudy (anonymous):

since Y is in the first quadrant the answer should be 6. just a sign error

OpenStudy (anonymous):

Got it thanks!

OpenStudy (anonymous):

no problem.